# references.bib

### v1

@@ 5.3 Colorectal Cancer
Table 4: Colorectal
: Tissue-specific model performance when applied to the SemiCOL challenge data. The model consist of organ-specific MIL head and frozen encoders: ImageNet, CTP, TC-Swin, TC-Conv. The table shows mean accuracy, F1 score, AUC, and the respective standard deviations over 5 different runs. Each model was trained on images from one site and evaluated on all images from the remaining sites. The best score is shown in bold. Clinic Model Acc F1 Auc LMU ImgNet 0.877 ±$\pm 0.05$ 0.875 ±$\pm 0.06$ 0.961 ±$\pm 0.01$ CTP 0.967 ±$\pm 0.01$ 0.964 ±$\pm 0.01$ 0.991 ±$\pm 0.004$ TC-Swin 0.948 ±$\pm 0.03$ 0.948 ±$\pm 0.03$ 0.987 ±$\pm 0.01$ TC-Conv 0.921 ±$\pm 0.02$ 0.919 ±$\pm 0.02$ 0.979 ±$\pm 0.01$ UBERN ImgNet 0.734 ±$\pm 0.09$ 0.709 ±$\pm 0.11$ 0.939 ±$\pm 0.02$ CTP 0.860 ±$\pm 0.08$ 0.854 ±$\pm 0.09$ 0.985 ±$\pm 0.01$ TC-Swin 0.943 ±$\pm 0.02$ 0.944 ±$\pm 0.02$ 0.987 ±$\pm 0.01$ TC-Conv 0.943 ±$\pm 0.01$ 0.943 ±$\pm 0.01$ 0.986 ±$\pm 0.01$ UKK ImgNet 0.890 ±$\pm 0.02$ 0.889 ±$\pm 0.02$ 0.960 ±$\pm 0.01$ CTP 0.962 ±$\pm 0.01$ 0.962 ±$\pm 0.01$ 0.993 ±$\pm 0.01$ TC-Swin 0.955 ±$\pm 0.01$ 0.953 ±$\pm 0.01$ 0.989 ±$\pm 0.01$ TC-Conv 0.925 ±$\pm 0.02$ 0.925 ±$\pm 0.01$ 0.978 ±$\pm 0.01$ WNS ImgNet 0.936 ±$\pm 0.02$ 0.926 ±$\pm 0.02$ 0.984 ±$\pm 0.004$ CTP 0.989 ±$\pm 0.01$ 0.989 ±$\pm 0.01$ 0.998 ±$\pm 0.001$ TC-Swin 0.967 ±$\pm 0.01$ 0.966 ±$\pm 0.01$ 0.993 ±$\pm 0.01$ TC-Conv 0.973 ±$\pm 0.01$ 0.973 ±$\pm 0.01$ 0.995 ±$\pm 0.01$ Average ImgNet 0.859 ±$\pm 0.09$ 0.852 ±$\pm 0.1$ 0.962 ±$\pm 0.02$ CTP 0.944 ±$\pm 0.06$ 0.943 ±$\pm 0.01$ $\mathbf{0.992}\pm 0.01$ TC-Swin $\mathbf{0.953}\pm 0.02$ $\mathbf{0.953}\pm 0.02$ 0.989 ±$\pm 0.01$ TC-Conv 0.941 ±$\pm 0.02$ 0.940 ±$\pm 0.02$ 0.984 ±$\pm 0.01$
@@ 5.4 Lung Cancer
As an additional experiment, the UNI [chen2024towards] encoder was compared to the other models using the ABMIL aggregation head and showed similar results to the other encoders. The detailed results can be found in B.1.
@@ 6 Discussion
Even though, performance generally is comparable, there are two differences between TC and CTP in organ specific performance. One is the difference on the prostate dataset. While a small amount of patches of this dataset were present during the pre-training of TC, the vast amount of patches used during testing were unseen. However, a lot of prostate patches were used during the pre-training of TC, as mentioned in section 3.1. This dominant representation of prostate tissue during pre-training could potentially lead to the better performance on this organ. This raises the question if organ-specific fine-tuning of TC encoders can lead to better performance on specific organs when considering cross-center evaluation and testing on unseen datasets. This question is further justified by the differences in organ specific performance between foundation models and external baselines.

### v2

@@ 5.3 Colorectal Cancer
Table 4: Colorectal: Tissue-specific model performance when applied to the SemiCOL challenge data. The model consist of organ-specific MIL head and frozen encoders: ImageNet, CTP, TC-Swin, TC-Conv. The table shows mean accuracy, F1 score, AUC, and the respective standard deviations over 5 different runs. Each model was trained on images from one site and evaluated on all images from the remaining sites. The best score is shown in bold.
Clinic Model Acc F1 Auc
LMU ImgNet 0.877 ±$\pm 0.05$ 0.875 ±$\pm 0.06$ 0.961 ±$\pm 0.01$
CTP 0.967 ±$\pm 0.01$ 0.964 ±$\pm 0.01$ 0.991 ±$\pm 0.004$
TC-Swin 0.948 ±$\pm 0.03$ 0.948 ±$\pm 0.03$ 0.987 ±$\pm 0.01$
TC-Conv 0.921 ±$\pm 0.02$ 0.919 ±$\pm 0.02$ 0.979 ±$\pm 0.01$
UBERN ImgNet 0.734 ±$\pm 0.09$ 0.709 ±$\pm 0.11$ 0.939 ±$\pm 0.02$
CTP 0.860 ±$\pm 0.08$ 0.854 ±$\pm 0.09$ 0.985 ±$\pm 0.01$
TC-Swin 0.943 ±$\pm 0.02$ 0.944 ±$\pm 0.02$ 0.987 ±$\pm 0.01$
TC-Conv 0.943 ±$\pm 0.01$ 0.943 ±$\pm 0.01$ 0.986 ±$\pm 0.01$
UKK ImgNet 0.890 ±$\pm 0.02$ 0.889 ±$\pm 0.02$ 0.960 ±$\pm 0.01$
CTP 0.962 ±$\pm 0.01$ 0.962 ±$\pm 0.01$ 0.993 ±$\pm 0.01$
TC-Swin 0.955 ±$\pm 0.01$ 0.953 ±$\pm 0.01$ 0.989 ±$\pm 0.01$
TC-Conv 0.925 ±$\pm 0.02$ 0.925 ±$\pm 0.01$ 0.978 ±$\pm 0.01$
WNS ImgNet 0.936 ±$\pm 0.02$ 0.926 ±$\pm 0.02$ 0.984 ±$\pm 0.004$
CTP 0.989 ±$\pm 0.01$ 0.989 ±$\pm 0.01$ 0.998 ±$\pm 0.001$
TC-Swin 0.967 ±$\pm 0.01$ 0.966 ±$\pm 0.01$ 0.993 ±$\pm 0.01$
TC-Conv 0.973 ±$\pm 0.01$ 0.973 ±$\pm 0.01$ 0.995 ±$\pm 0.01$
Average ImgNet 0.859 ±$\pm 0.09$ 0.852 ±$\pm 0.1$ 0.962 ±$\pm 0.02$
CTP 0.944 ±$\pm 0.06$ 0.943 ±$\pm 0.01$ $\mathbf{0.992}\pm 0.01$
TC-Swin $\mathbf{0.953}\pm 0.02$ $\mathbf{0.953}\pm 0.02$ 0.989 ±$\pm 0.01$
TC-Conv 0.941 ±$\pm 0.02$ 0.940 ±$\pm 0.02$ 0.984 ±$\pm 0.01$
@@ 5.4 Lung Cancer
As an additional experiment, the UNI [chen_towards_2024] encoder was compared to the other models using the ABMIL aggregation head and showed similar results to the other encoders. The detailed results can be found in B.1.
@@ 6 Discussion
Even though, performance generally is comparable, there are two differences between TC and CTP in organ specific performance. One is the difference on the prostate dataset. While a small amount of patches of this dataset were present during the pre-training of original TC encoder, the vast amount of patches used during testing were unseen. Overall, a lot of prostate patches were used during the pre-training of TC, as mentioned in section 3.1. This large proportion of prostate tissue during pre-training may have led to the better performance on this organ and it likely prevented a larger reduction of the encoder’s performance when omitting the small amount of PANDA data during pre-training. This raises the question if organ-specific fine-tuning of TC encoders can lead to better performance on specific organs when considering cross-center evaluation and testing on unseen datasets. This question is further justified by the differences in organ specific performance between foundation models and external baselines.